# RL Zero: Zero-Shot Language to Behaviors without any supervision

### v1

@@ 3 Preliminaries
Multimodal Video-Foundation Models (ViFMs) and In-Domain Video Generation: Multimodal ViFMs (Wang et al., 2022; 2024a; Tong et al., 2022) facilitate the understanding of video data in a shared representation space of modalities such as text or audio. Recent works in video generation (Kondratyuk et al., 2023; Blattmann et al., 2023) have used this capability to condition the model of various input modalities that are first converted to this shared representation space. Notably, these models can utilize text prompts to guide content, style, and motion, or employ an image as the initial frame of a subsequent video sequence. For this work, we use off-the-shelf video generation models $V\!M$ that predict a sequence of video frames $\{i_{1},i_{2},...i_{n}\}$ given a task specified in natural language $l$ by first projecting the language in a common embedding space across modalities. Formally, $V\!M:l\rightarrow\{i_{1},i_{2},...i_{n}\}$.
Zero-shot RL using Successor Measures: Successor Measures (Blier et al., 2021a) have been recently studied (Touati & Ollivier, 2021; Agarwal et al., 2024) for their ability to obtain zero-shot RL policies for any given reward function. Mathematically, successor measures define the measure over future states visited as $M^{\pi}$,
$M^{\pi}(s,a,X)=\mathbb{E}_{\pi}[\sum_{t\geq 0}\gamma^{t}p^{\pi}(s_{t+1}\in X|s,a)]~{}~{}\forall X\subset\mathcal{S}.$ (1)
Successor Features (Dayan, 1993) are the discounted sum of state features under a policy $\pi$ and are given as,
$\psi^{\pi}(s,a)=\mathbb{E}_{\pi}[\sum_{t\geq 0}\gamma^{t}\phi(s_{t})],$ (2)
where $\phi:S\rightarrow\mathbb{R}^{d}$ are state features spanning reward functions using $r(s)=\phi^{T}z$, and $z$ are the linear weights. It has been shown that learning successor measures can be used to jointly obtain successor features along with these state features.
Lemma 1.
(Theorem 13 of Touati et al. (2023)) For an offline dataset with density $\rho$, if the successor measure is represented as $M^{\pi}(s,a,s^{+})=\psi^{\pi}(s,a)\varphi(s^{+})\rho(s^{+})$, then $\psi$ is the successor feature $\psi^{\pi}(s,a)$ for state feature $\phi(s)=\varphi(s)^{T}(\mathbb{E}_{\rho}(\varphi\varphi^{T}))^{-1}$.
The zero-shot RL pipeline thus works in two steps: pretraining and zero-shot inference. To represent a distribution over all policies, the policy can be conditioned on $z$, so the successor feature can be represented as $\psi(s,a,z)$.
Step 1 – Pretraining: Obtain successor features and state features jointly by learning $M^{\pi_{z}}$ parameterized as $\psi(s,a,z)\varphi(s^{+})$,
$M^{\pi_{z}}(s,a,X)=\int_{s^{+}\in X}\psi(s,a,z)^{\top}\varphi(s^{+})\rho(s^{+})~{}~{}\forall z,X\subset\mathcal{S}$ (3)
Furthermore, the policy $\pi_{z}$ can be co-trained to be the near-optimal policy under reward function $r(s)=\phi(s)^{T}z$, thus defining a mapping from reward to optimal policies. This is implemented by setting $\pi_{z}$ to output greedy actions with maximum $Q$-value under the reward function $r$.

### v2

@@ 3 Preliminaries
Multimodal Video-Foundation Models (ViFMs) and In-Domain Video Generation: Multimodal ViFMs [78, 79, 75] enable the understanding of video data in a shared representation space of other modalities such as text or audio. These shared representations can be used to condition video generation on different input modalities [36, 7]. Notably, these models can utilize text prompts to guide content, style, and motion, or employ an image as the initial frame for a subsequent video sequence. For this work, we use video generation models $V\!M$ that generate a sequence of video frames $\{i_{1},i_{2},...i_{n}\}$ given a task specified in natural language $l$ by first converting the language prompt to a common embedding space across modalities; formally, $V\!M:l\rightarrow\{i_{1},i_{2},...i_{n}\}$.
Imitation Learning through Distribution Matching: We consider a learning agent in a Markov Decision Process (MDP) [58, 71] which is defined as a tuple: $\mathcal{M}=({\mathcal{S}},{\mathcal{A}},p,r,\gamma,d_{0})$ where ${\mathcal{S}}$ and ${\mathcal{A}}$ denote the state and action spaces respectively, $p$ denotes the transition function with $p(s^{\prime}|s,a)$ indicating the probability of transitioning from $s$ to $s^{\prime}$ taking action $a$; $r$ denotes the reward function, $\gamma\in(0,1)$ specifies the discount factor and $d_{0}$ denotes the initial state distribution. The reinforcement learning objective is to obtain a policy $\pi:{\mathcal{S}}\rightarrow\Delta({\mathcal{A}})$ that maximizes expected return: $\mathbb{E}_{\pi}[{\sum_{t=0}^{\infty}\gamma^{t}r(s_{t})}]$, where we use $\mathbb{E}_{\pi}$ to denote the expectation under the distribution induced by $a_{t}\sim\pi(\cdot|s_{t}),s_{t+1}\sim p(\cdot|s_{t},a_{t})$ and $\Delta({\mathcal{A}})$ denotes a probability simplex supported over ${\mathcal{A}}$.
An imitation learning agent does not have access to the reward function, $R$, but has access to an "expert" trajectory (or a set of "expert" trajectories) from a policy that maximizes the reward function. Distribution matching objectives [25, 52] for imitation learning have been commonly used in some recent work [24, 66]. The objective in these methods is $\min_{\pi}{\mathcal{D}}(\rho^{\pi},\rho^{E})$, where $\rho^{\pi}$ is the visitation distribution of the policy $\pi$ (defined by the probability of being in state $s$ starting from the initial state distribution $s_{0}$ and following the policy $\pi$), $\rho^{E}$ is the visitation distribution exhibited by the "expert" trajectory and ${\mathcal{D}}$ is a function to compare the closeness of the distributions. $f$-Divergences are commonly used as a measure of distance between distributions.
Successor Measure based Unsupervised RL: Successor Measure [8] based learning has been recently studied [76, 2] as an unsupervised RL objective for its ability to describe long-term behavior of the policy in the environment. Mathematically, successor measures define the measure over future states visited as $M^{\pi}$,
$M^{\pi}(s,a,X)=\mathbb{E}_{\pi}\left[\sum_{t\geq 0}\gamma^{t}p^{\pi}(s_{t+1}\in X|s,a)\right]~{}~{}\forall X\subset\mathcal{S}.$ (1)
Representing the successor measure for any policy $\pi$ as $\psi^{\pi}(s,a)^{T}\varphi(s^{+})$ and trained on a dataset $d^{O}$ of offline interactions of the form $\{s,a,s^{\prime}\}$, these methods facilitate extraction of a state-representation $\varphi(s)$ that is suitable for RL. Then, policies $\pi_{z}$(where the policies are represented using latents $z$) are learned offline that are near-optimal for a reward function defined in the span of learned state-features $r_{z}(s)=\varphi(s)\cdot z$. At test time, the policy for any given reward function ($r(s)$) can then be obtained with such a BFM (zero-shot, with no additional experiential data or gradient updates) using a closed-form solution to the following linear regression:
$z_{R}=\min_{z}\mathbb{E}_{d^{O}}[(r(s)-\varphi(s)\cdot z)]^{2}\implies z_{R}=\mathbb{E}_{d^{O}}[\varphi(s)\cdot\varphi(s)^{T}]^{-1}\mathbb{E}_{d^{O}}[\varphi(s)\cdot r(s)]$ (2)
For this work, we use ($\varphi(s),\pi_{z}$) to represent a BFM trained with successor-measure based unsupervised RL.
4 RLZero: Zero-Shot Prompt to Policy
RLZero uses components trained without any explicit in-domain supervision to map language to behaviors. For each domain, we consider a dataset $d^{O}$ of reward-free interactions {$s,a,s^{\prime}$} and a BFM ($\varphi(s),\pi_{z}$) pre-trained on $d^{O}$. In the following sections, we describe the steps involved in detail. First, we present how an imagined trajectory is generated from a prompt. Then, we discuss how this imagined trajectory is projected to real observations of an agent. Finally, we describe the zero-shot procedure for inferring a policy that matches the behavior in the imagined trajectory. The latter two steps can also be used to directly infer policy when instead of language prompt, a cross-embodied video demonstration is provided.
4.1 Imagine: Unsupervised Generative Text-Conditioned Video Modeling
Grounding language to tasks has historically [26, 33, 54] required costly annotation labels that map language to task examples specified through image or state trajectories. Large video-language foundation models (ViFMs) help lift that requirement by training on vast amounts of internet videos, thus giving us a rich prior of grounding language commands to videos. We leverage the unsupervised video modeling approach from GenRL [49]. In the absence of any in-domain text labels, the approach uses an off-the-shelf video-language encoder with video encoder $f_{v}$ and language encoder $f_{l}$ trained on internet scale data to encode a sequence of observation frames ($o_{1:T}$) to a embedding space shared with its language description. A generative video model conditions on this latent embedding to generate a sequence of image observations ($i_{1:T})$. The video model ($V\!M$) is trained to reconstruct the original observation frames ($o_{1:T}$) and at test time can simply be queried with a language instruction. Similar to GenRL, we use InternVideo2 as the video-language encoder and use a simplified GRU architecture for video generation. We refer the readers to GenRL for more details on training the video generation model and accounting for multimodality gap between video-text encodings. Training the video generation model does not require text labels mapping language to tasks and is fully unsupervised. Thus during inference, given a language instruction $e^{l}$, we obtain a sequence of frames $V\!M(f_{l}(e^{l}))=(i_{1},i_{2}...i_{T})$ that represents an imagination of what the task looks like in the environment domain. Figure 6 shows an example of what these imaginings look like using the trained video generation model. Given the generated video of text description, GenRL learns an action-conditional world model to train a new policy for each text-instruction with model-based RL. This makes policy inference for a variety of tasks costly, unsafe, and memory-inefficient. In contrast, with RLZero we present a way to lift this limitation with unsupervised RL to directly infer a policy given generated video in the sections below.
4.2 Project: Grounding to Agent’s Observation Space
The imaginings produced by $V\!M$ can be noisy, unrealizable, and not exactly representative of the domain. We propose to use semantic-similarity based retrieval to obtain nearest frames in the dataset of the agent’s prior environmental interactions $d^{O}$ to project individual frames of imagined trajectories to real observations as shown in Fig 2.
Figure 2: Grounding Imagination in Real Observations: We use nearest image retrieval defined by cosine similarity in the embedding space to output a real observation from the dataset that matches the imagined frame.
Semantic matching also allows us the flexibility to replace imaginings with a video demonstration of a task by a different embodiment agent in a different domain (e.g. as we will demonstrate in Section 5.2). In this work, we use an off-the-shelf vision-language embedding approach for retrieval, SigLIP [89], to map both the imagined frame and agent observation to the same latent embedding space, which has been pre-trained for similarity matching on an internet-scale dataset. As a single image observation is not informative of key state variables such as velocity or acceleration, we use the frame stacking technique common in RL [50] when finding the nearest observation from the dataset corresponding to any imagined frame. We use an encoding function $\mathcal{E}:\mathcal{I}\rightarrow\mathcal{Z}$ to individually map a sequence of images to shared image-text embedding space.:
$o_{t}=\operatorname*{arg\,max}_{o_{t}}\frac{\mathcal{E}(o_{t-k:t})\cdot\mathcal{E}(i_{t-k:t})}{\|\mathcal{E}(o_{t-k:t})\|\|\mathcal{E}(i_{t-k:t})\|}~{}~{}~{}\forall t\in[T].$ (3)
Above we use $k$ previous frames (usually set to 3) to identify state variables in order to improve matching accuracy. Using the offline interaction dataset, we find corresponding proprioceptive states in addition to the real observation that we will subsequently use as a target sequence to imitate. We rely on proprioceptive states for the subsequent unsupervised RL phase, but in general BFMs may be trained with image observations if the state information is unavailable. While this step allows us to ground imagination to a sequence of real observations that the video expects the agent to follow, action inference still remains a challenge as the sequence might not be dynamically feasible or be available as a trajectory in the dataset to behavior clone.
4.3 Imitate: Zero-Shot Distribution Matching with Unsupervised RL
The resulting projected sequence of states might be dynamically infeasible and out of distribution from the trajectories in the offline dataset. Our key insight is to leverage unsupervised RL techniques to discover new skills from an offline dataset and use state-marginal matching to find a policy that induces the closest dynamically feasible sequence imitating the imagination. This section demonstrates that after pretraining a successor-measure based unsupervised RL agent with reward-free transitions, an approximate closed-form solution to state marginal matching exists without requiring any further gradient steps or interaction with the environment.
We take the distribution matching perspective of imitation learning [28, 25] and find a policy that matches the state marginal distributions of the grounded imagined trajectories (expert). Given a dataset of offline environmental interactions ($d^{O}$ with distribution $\rho$), we pretrain a successor-measures based unsupervised RL agent [2, 77] ($\varphi(s),\pi_{z}$) that provides us with a mapping from reward function $R$ to the corresponding $z_{R}$ that induces a near-optimal policy $\pi_{z_{R}}$ as shown in Eq. 2. For the case of state-marginal matching, finding the optimal policy reduces to finding the optimal latent $z$ minimizing the following distribution matching objective:
$z_{imit}=\operatorname*{arg\,min}_{z}{\mathcal{D}}(\rho^{\pi_{z}}(s),\rho^{E}(s)),$ (4)
where $\rho^{E}$ and $\rho^{\pi_{z}}$ are state visitation distribution of the "expert" imagined trajectories and of the policy $\pi^{z}$ respectively and ${\mathcal{D}}$ can be chosen to be mean-squared error, $f$-divergence, Integral Probability Metrics (IPM), etc. In general, minimizing the distance via gradient descent can provide a solution $z_{imit}$ to distribution matching. For the special case of KL divergence, Theorem 1 shows that $z_{imit}$ can be obtained using a learned distribution ratio between expert and offline interaction dataset $\rho^{E}/\rho$.
Theorem 1.
Define $J(\pi,r)$ to be the expected return of a policy $\pi$ under reward $r$. For an offline dataset $d^{O}$ with density $\rho$, a learned log distribution ratio: $\nu(s)=\log(\frac{\rho^{E}(s)}{\rho(s)})$, $D_{KL}(\rho^{\pi},\rho^{E})\leq-J(\pi,r^{imit})+D_{KL}(\rho^{\pi}(s,a),\rho(s,a))$ where $r^{imit}(s)=\nu(s)~{}\forall s$. The corresponding $z_{imit}$ minimizing the upper bound is given by $z_{imit}=\mathbb{E}_{\rho}[{r^{imit}(s)}\varphi(s)]=\mathbb{E}_{\rho^{E}}[\frac{\nu(s)}{e^{\nu(s)}}\varphi(s)]$ where $\varphi$ denotes state features learned by the BFM.
Thus, with the reward function $r^{imit}(s)=\nu(s)$, we can use the solution of $z_{imit}=\mathbb{E}_{\rho}{[\varphi(s)r^{imit}(s)]}$ to retrieve the policy that mimics the grounded imagined behavior. However this reward function still requires learning a discriminator to obtain the distribution ratio for each task, which can lead to instabilities. A heuristic yet performant alternative is to use a shaped reward function $r(s)=e^{\nu(s)}$, similar to Pirotta et al. [57], which allows closed-form inference ($z_{imit}=\mathbb{E}_{\rho}{[e^{\nu(s)}\varphi(s)]}=\mathbb{E}_{\rho}{[\frac{\rho^{E}(s)}{\rho(s)}\varphi(s)]}=\mathbb{E}_{\rho^{E}}{[\varphi(s)]}$) without learning a discriminator, any gradient steps or interaction with the environment. We compare both approaches in Appendix C.2. The performance for both these methods are almost identical and we defer to the latter one (gradient-free inference approach) in all our experiments. The complete algorithm for RLZero can be found in Algorithm 1.
5 Experiments
Our experiments seek to understand the quality of behaviors that RLZero is able to produce in two settings (a) Language to Behavior, and (b) Cross-embodied Video to Behavior Generation. We note that none of the experiments below assume in-domain supervision such as annotations of trajectories in the environment with their task label, or expert demonstrations corresponding to the specified tasks either for training the video generation model or training the behavioral foundation model.
{NiceTabular}
cc|cc|cc|c|c Image-language reward (VLM-RM) Video-language reward GenRLRLZero
IQL TD3 (Base Model) TD3 IQL model-based
Walker Lying Down 2/5 (307.04$\pm$52.60) — (116.97$\pm$69.70) 2/5 (120.48$\pm$50.47) 5/5 (419.05$\pm$281.56)4/5 (199.66$\pm$30.64) 5/5 (524.34$\pm$31.78)
Walk like a human 1/5 (95.05$\pm$78.08) — (30.34$\pm$34.58) 3/5 (49.49$\pm$45.73) 4/5 (146.86$\pm$54.40) 5/5 (632.09$\pm$36.31)5/5 (704.68$\pm$47.16)
Run like a human 5/5 (50.93$\pm$5.05) — (61.76$\pm$52.49) 1/5 (100.24$\pm$89.76) 2/5 (286.14$\pm$63.91) 5/5 (316.21$\pm$82.61) 5/5 (475.49$\pm$52.37)
Do lunges 4/5 (130.98$\pm$109.87) — (67.24$\pm$24.31) 2/5 (48.71$\pm$12.06) 3/5 (217.78$\pm$147.57) 5/5 (484.68$\pm$50.60) 5/5 (377.53$\pm$30.14)
Cartwheel 4/5 (320.03$\pm$228.81) — (167.04$\pm$48.12) 3/5 (151.16$\pm$67.95) 4/5 (175.45$\pm$50.87) 5/5 (252.68$\pm$33.76) 4/5 (254.89$\pm$12.74)
Strut like a horse 5/5 (110.08$\pm$50.24) — (164.61$\pm$110.45) 1/5 (103.45$\pm$69.48) 3/5 (81.14$\pm$25.01) 5/5 (361.45$\pm$103.07) 5/5 (445.66$\pm$131.36)
Crawl like a worm 4/5 (116.48$\pm$72.99) — (94.54$\pm$16.18) 1/5 (84.43$\pm$36.97) 2/5 (127.14$\pm$10.11) 0/5 (127.48$\pm$24.81) 3/5 (143.86$\pm$23.09)
Quadruped Cartwheel 1/5 (23.82$\pm$8.38) — (6.81$\pm$3.24) 3/5 (9.45$\pm$7.43) 1/5 (14.07$\pm$5.33) 4/5 (15.6$\pm$4.07)4/5 (30.88$\pm$0.95)
Dance 5/5 (13.89$\pm$6.19) — (8.19$\pm$4.74) 3/5 (6.24$\pm$2.34) 1/5 (11.79$\pm$4.28) 4/5 (15.11$\pm$7.71) 5/5 (24.29$\pm$3.29)
Walk using three legs 2/5 (15.98$\pm$9.61) — (6.70$\pm$3.61) 2/5 (7.96$\pm$4.97) 3/5 (7.46$\pm$1.75) 4/5 (24.70$\pm$6.56) 5/5 (27.29$\pm$7.29)
Balancing on two legs 2/5 (13.92$\pm$5.52) — (6.11$\pm$3.95) 2/5 (6.69$\pm$8.19) 2/5 (5.61$\pm$3.05) 5/5 (7.20$\pm$3.78) 5/5 (22.50$\pm$2.15)
Lie still 1/5 (82.93$\pm$77.13) — (9.51$\pm$7.54) 3/5 (6.88$\pm$8.43) 2/5 (67.96$\pm$34.91) 3/5 (135.26$\pm$57.41) 2/5 (100.22$\pm$52.45)
Handstand 2/5 (16.13$\pm$11.66) — (4.18$\pm$0.88) 4/5 (4.08$\pm$2.74) 2/5 (6.80$\pm$2.85) 4/5 (15.47$\pm$11.76) 3/5 (50.65$\pm$2.92)
Cheetah
Lie down 3/5 (219.35$\pm$58.77) — (203.99$\pm$16.39) 2/5 (255.03$\pm$49.28) 3/5 (360.76$\pm$43.80) 3/5 (468.52$\pm$17.45) 2/5 (202.67$\pm$34.75)
Bunny hop 3/5 (218.44$\pm$34.56) — (181.74$\pm$35.52) 1/5 (192.78$\pm$53.53) 3/5 (129.26$\pm$1.37) 4/5 (220.62$\pm$36.99)5/5 (224.70$\pm$6.46)
Jump high 3/5 (172.10$\pm$39.02) — (184.41$\pm$49.01) 0/5 (183.36$\pm$73.40) 5/5 (148.82$\pm$27.88) 5/5 (267.54$\pm$28.17)5/5 (232.23$\pm$41.65)
Jump on back legs and backflip 3/5 (175.41$\pm$38.86) — (197.62$\pm$42.52) 0/5 (169.67$\pm$50.78) 2/5 (131.14$\pm$15.24) 5/5 (293.28$\pm$71.71)5/5 (326.20$\pm$57.45)
Quadruped walk 3/5 (379.34$\pm$21.07) — (193.64$\pm$20.30) 3/5 (187.24$\pm$46.67) 3/5 (188.58$\pm$20.95) 2/5 (318.10$\pm$37.53)4/5 (388.02$\pm$37.76)
Stand in place like a dog 4/5 (478.85$\pm$43.69) — (282.46$\pm$94.73) 3/5 (238.58$\pm$88.26) 0/5 (169.39$\pm$14.04)2/5 (238.25$\pm$47.83) 3/5 (469.05$\pm$31.62)
Stickman Lie down stable 2/5 (201.03$\pm$82.97) — (13.21$\pm$9.27) 4/5 (30.86$\pm$4.60) 1/5 (28.89$\pm$3.49) 5/5 (686.56$\pm$386.66) 4/5 (841.48$\pm$226.24)
Lunges 0/5 (63.46$\pm$36.27) — (249.73$\pm$13.81) 2/5(48.69$\pm$25.50) 0/5 (38.77$\pm$3.69) 4/5 (244.82$\pm$58.80) 5/5 (191.41$\pm$61.51)
Praying 1/5 (50.84$\pm$22.27) — (39.39$\pm$42.79) 0/5 (49.13$\pm$32.51) 0/5 (40.76$\pm$9.17) 3/5 (192.75$\pm$42.20)4/5 (147.74$\pm$54.07)
Headstand 2/5 (20.14$\pm$12.24) — (14.54$\pm$7.13) 2/5 (48.11$\pm$42.51) 1/5 (13.84$\pm$6.02) 4/5 (71.75$\pm$32.77)4/5 (71.87$\pm$6.28)
Punch 2/5 (73.42$\pm$19.00) — (50.23$\pm$38.08)3/5 (74.73$\pm$7.01) 4/5 (85.76$\pm$22.96) 5/5 (181.08$\pm$69.58) 4/5 (216.44$\pm$37.81)
Plank 0/5 (374.70$\pm$361.27) — (507.46$\pm$289.41) 0/5 (16.19$\pm$5.70) 0/5 (66.62$\pm$53.43) 1/5 (391.04$\pm$41.15)3/5 (883.60$\pm$58.00)